Given the function f(x)=6x+7
f(3)=
f(-4)=
f(6.2)=

Answer by stanbon(75887)

Given the function f(x)=6x+7
f(3)= 6*3 + 7 = 18+7 = 25
-----------------------------
f(-4)= 6*-4 + 7 = -24 +7 = -17
-----------------------------
f(6.2) = 6*6.2 + 7 = 44.2
=================================
Cheers,
Stan H.